A material has conductivity of 10-2 mho/m and a relative permittivity of 4. The frequency at which conduction current in the medium is equal to displacement current is


1
Expert's answer
2021-01-14T10:38:48-0500

Answer

Frequency can be given as

"f=\frac{\sigma }{2\pi \epsilon}=\frac{10^{-2}}{2\pi \times4\times8.85\times10^{-12}}=45MHz"